OpenStudy (anonymous):

The equation s=p(1+r)^t can be used to model a person’s future income. S = future salary, p = current salary, r = rate of increase(% in decimal form), and t = time in years. Bob’s salary increases by 5% each year. If he starts at $35,000 per year, what will his salary be in 5 years?

jimthompson5910 (jim_thompson5910):

plug in p = 35000, r = 0.05 and t = 5

jimthompson5910 (jim_thompson5910):

s = p(1+r)^t s = 35000(1+0.05)^5 s = ???

OpenStudy (anonymous):

Thank you I got it

jimthompson5910 (jim_thompson5910):

yw